OpenStudy (anonymous):

write in vertex form and graph y=x^2+4x+8

jigglypuff314 (jigglypuff314):

vertex form for a parabola: y = a(x - h)^2 + k (h, k) being the vertex x^2 + 4x + 4 = (x + 2)^2 so how would you get x^2 + 4x + 8 like that? try x^2 + 4x + 4 + 4 so now you have (x + 2)^2 + 4